OpenStudy (anonymous):

what is the quotient of (2.965*10^7) / (5*10^3) in scientific notation

OpenStudy (phi):

this one is a little trickier break it into \[ \frac{2.965}{5} \cdot \frac{10^7}{10^3} \] for the 10^7 divided by 10^3 you subtract exponents. top exponent minus bottom exponent. can you do that ?

OpenStudy (anonymous):

yup

OpenStudy (phi):

and divide 2.965 by 5

OpenStudy (phi):

but you are not quite done yet. so far, what do you have ?

OpenStudy (anonymous):

alright got it but what in this in a recent year, the world production of wheat was about 5.2*10^8 metric tons,canada produced 1/25 of this wheat.About how much wheat did canada produce?

OpenStudy (phi):

first, what did you get for the question you posted ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

answer choice are 20.8*10^8 2.08*10^7 2.08*10^6 20.8*10^7